OpenStudy (anonymous):

Please, some one intelligent here, could help to know, how this two triangles are similar??

OpenStudy (anonymous):

the question?

OpenStudy (anonymous):

 to prove if this two triangle are similar .

OpenStudy (anonymous):

maybe I could use ASA to prove?? I not pretty sure

OpenStudy (anonymous):

AA , if two angles on triangles equal to two angles of another triangle,they are similar :)

OpenStudy (anonymous):

we don't have ASA! AA,SS

OpenStudy (debbieg):

If 2 angles are equal then the 3rd must be equal..... but do we know anything about the sides? can have infinitely many triangles with the same 3 angle measures.

OpenStudy (anonymous):

exactly

OpenStudy (debbieg):

oh right.... we only want them to be similar, not congruent.... :) so, yeah, you just need to notice that if the 2 angles are the same, then the 3rd one is also.

OpenStudy (debbieg):

Similar just means that the angles are all the same. They are the same "shape", not necessarily the same SIZE. :)

OpenStudy (anonymous):

yes you are right thanks

OpenStudy (anonymous):

and when, \[\frac{ S _{1} }{ S _{2} } = \frac{ S _{3} }{ S _{4} } = \frac{ S _{5} }{ S _{6} }\] they are similar :)

OpenStudy (anonymous):

got it?

OpenStudy (anonymous):

yes thanks a lot guys

OpenStudy (anonymous):

you're welcome :)